The American Diabetes Association recommends a blood glucose reading of less than 130 for those with Type 2 diabetes. Blood glucose measures the amount of sugar in the blood. Below are the readings for February for a person recently diagnosed with Type 2 diabetes.

119 106 98 115 126 114 119 109 119 114
97 111 124 120 109 107 109 119 97 123
125 114 97 106 122 111 113 107

Click here for the Excel Data File

  

a. What is the arithmetic mean glucose reading? (Round your answer to 2 decimal places.)

b. What is the median glucose reading? (Round your answer to 1 decimal place.)

c. What is the modal glucose reading?
Answer #1

Arithmetic Mean: Σxi / n = 3150 / 28 = 112.5

Median = If number of data points are even, then median is the mean of the middlemost numbers.

Median = 113+114/2 = 113.5

Modal value: Modal value is the most repeated observation in the data set.

Modal value = 119